# A Unified Approach to Emotion Detection and Task-Oriented Dialogue Modeling

### v1

@@ Abstract
In current text-based task-oriented dialogue (TOD) systems, user emotion detection (ED) is often overlooked or is typically treated as a separate and independent task, requiring additional training. In contrast, our work demonstrates that seamlessly unifying ED and TOD modeling brings about mutual benefits, and is therefore an alternative to be considered. Our method consists in augmenting SimpleToD, an end-to-end TOD system, by extending belief state tracking to include ED, relying on a single language model. We evaluate our approach using GPT-2 and Llama-2 on the EmoWOZ benchmark, a version of MultiWOZ annotated with emotions. Our results reveal a general increase in performance for ED and task results. Our findings also indicate that user emotions provide useful contextual conditioning for system responses, and can be leveraged to further refine responses in terms of empathy.
@@ 1 Introduction
Emotional user utterances frequently occur during interactions with dialogue systems reeves1996media (19). In open-domain dialogues, users directly share personal and emotional experiences li-etal-2017-dailydialog (10). In task-oriented dialogues (TODs), emotions are closely related to task progression. They usually become apparent as the task unfolds, and are contingent on the user’s expectations being met feng-etal-2022-emowoz (5). Detecting these emotions explicitly in either scenario is beneficial for several reasons: it can help with reviewing chat logs after the exchange concludes or with adjusting responses during the exchange to better align with the user’s emotional state rashkin-etal-2019-towards (18). In TODs, these more empathetic responses have demonstrated their effectiveness in compensating for system errors i-feel-you-2013 (12), creating the impression of a more capable system.
Existing text-based approaches to ED in TODs either require a dedicated, specifically trained component feng-etal-2022-emowoz (5, 6) or assume implicit ED. This assumption is due to systems being commonly trained to replicate human expert responses xu2017new (25, 7) which inherently convey empathy, as needed, and therefore implicit ED. In contrast, we propose an approach that eliminates the need for training an additional component while explicitly modeling user emotions, applicable in settings where emotion annotations are available. Our method involves predicting user emotions within a sequence of task-related components, treating ED as an extension of belief state tracking. Our approach is end-to-end and relies on a single language model, optimized in a unified fashion. Specifically, we build upon SimpleTOD hosseini-asl_simple_2020 (8), a popular end-to-end TOD approach trained on GPT-2 radford2019gpt2 (17) (Sec. 2). Recognizing the success of large language models (LLMs) brown2020language (1), we also fine-tune a state-of-the-art LLM, Llama-2 touvron2023llama2 (22), using LoRA hu2021lora (9), a parameter-efficient fine-tuning technique (Sec. 2).
Moreover, we show that when considering Llama-2, ED predictions can be readily used to more explicitly condition and refine system responses. Capitalizing on the fact that LoRA weights are optimized on top of Llama-2’s fixed parameters, it is possible to dynamically disable the LoRA weights after generating TOD predictions to harness the LLM’s in-context learning ability. We refine responses without further training, using a chain-of-thought prompting approach wei2022chain (23) (Sec. 2).
We evaluate our method on the text-based EmoWOZ benchmark feng-etal-2022-emowoz (5), the only benchmark to provide emotion annotations on MultiWOZ budzianowski-etal-2018-multiwoz (2). The authors propose several competitive ED-only baselines, with ContextBERT feng-etal-2022-emowoz (5) being the most successful and taking the entire dialogue history as input.
@@ 2 Method and Experimental Setup
This is the only text-based dataset offering emotion annotations for a large and popular corpus of 10,000 TODs. It focuses on 7 user emotions: neutral, excited, fearful, satisfied, dissatisfied, apologetic, and abusive. These follow the OCC model (Ortony, Clore and Collins) ortony_clore_collins_1988 (15) and are based on 3 possible emotion elicitors which include the operator (or system), the user, and events. This emotion schema is highly specific to TODs and makes it possible to distinguish between emotions caused by the system and those provoked by external factors, independent of the system’s performance.
@@
Examples of user utterances and their associated emotion label can be found in table 5 and Appendix 4. Dataset statistics are shown in Table 1. Certain emotions are notably rare and make this a challenging benchmark for ED in TODs.
@@
Our approach builds upon SimpleToD hosseini-asl_simple_2020 (8), a popular end-to-end strategy that employs a single language model, similar to other recent end-to-end TOD systems ham-etal-2020-end (7, 26). In training, the model is exposed to instances where Context (all preceding turns $C_{t}=[U_{0},S_{0},...,U_{t}]$), Belief State ($B_{t}$), Dialogue Acts ($A_{t}$) and System Response ($S_{t}$) are concatenated into a single text sequence. To ensure database adaptability, responses are additionally delexicalized. During inference, only the Context $C_{t}$ is passed as input, and the model is expected to generate the rest of the sequence. Moreover, the predicted belief state is used to query the database to retrieve values with which to replace delexicalized placeholders. This pipeline can be readily augmented, and previous studies have in fact done so with snippets of knowledge chen-etal-2022-ketod (3) and chitchat sun-etal-2021-adding (21).
@@
The models used for our experiments include GPT-$2$https://huggingface.co/gpt2 radford2019gpt2 (17), typically used with SimpleToD, and the more recent, state-of-the-art Llama-2-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hftouvron2023llama2 (22). This is an open-source LLM trained on data from publicly accessible sources which can undergo further fine-tuning or be prompted directly with instructions and/or few-shot examples. To adapt models to our task, we fully fine-tune GPT-2 and employ LoRA fine-tuning for Llama-2, preserving a significant portion of the pre-trained weights. This approach enhances regularization, generally leading to improved generalization compared to a fully fine-tuned alternative. Training details can be found in Appendix 2.
When refining responses, we focus on the PREV model’s predictions. Once these have been generated, we can disable the LoRA weights and, with no additional training, prompt the model using a few-shot, chain-of-thought wei2022chain (23) approach (REFINE). This consists in creating demonstrations which include a reasoning step to prime the LLM’s output. Each few-shot demonstration consists in a) the dialogue Context, b) the emotion to be predicted, c) the response to be generated, d) a chain-of-thought reasoning step suggesting which behavior to display. Finally, it also contains e) an emotion-aware snippet that can be prepended to the previously generated response. Compared with refining the entire response, prompting the model to produce a prependable snippet mitigates task-related errors due to substantial modifications made by the LLM. We incorporate specific instructions into the prompt and construct one exemplar per emotion, picking examples from the training set. The demonstration snippets we design are meant to convey remorse, reassurance, appeasement, or enthusiasm, depending on the user’s emotion. Appendix 3 shows an illustrative example. When generating snippets, we filter those which are overly similar to the generated response to avoid redundancy, based on Levenshtein distance. Empirically, we find that filtering out snippets with a similarity ratio $\geq$ 50% is effective.
@@
For ED, we adhere to the EmoWOZ metrics, presenting F1 scores for each emotion alongside the macro and weighted averages excluding the neutral class. For end-to-end TOD, we employ the standard MultiWOZ metrics. The joint goal accuracy (JGA) reflects the proportion of turns where the predicted user constraints exactly match the reference ones. The inform rate evaluates the system’s capacity to provide the right type of entities from the database, given the user’s constraints, and the success rate assesses how effectively the system delivers requested attributes like phone numbers or booking references. For a more comprehensive understanding of these metrics, we direct readers to the MultiWOZ paper budzianowski-etal-2018-multiwoz (2). For response quality, we use the BLEU papineni-etal-2002-bleu (16) score. For response diversity, we analyze the Conditional Bigram Entropy (CBE) and the count of unique trigrams to gauge the richness of vocabulary and phrasing nekvinda-dusek-2021-shades (14).
@@ 3 Results and Discussion
For Llama-2, the PREV variant shows a clearer benefit. This model performs best overall, except for the dissatisfied emotion. By examining user utterances with this particular emotion, we find that they appear highly contextual and generally lack explicit signs of dissatisfaction, a common trait when identifying emotions from text only multimodal_emo_rec (20). The regularization from LoRA in this case is a drawback compared to full fine-tuning, as annotations for this emotion are dataset-specific. However, for less common emotions characterized by more explicit expressions, such as fearful and abusive, the use of an LLM with LoRA proves advantageous. Nonetheless, opportunities for further improvements remain, particularly in the recognition of abusive and dissatisfied, two crucial emotional states.
@@
On the other hand, the EMO variant does not offer as much signal, yielding results similar to the baseline in each case. Compared to GPT-2, Llama-2 responses exhibit greater diversity, as indicated by the CBE score and the unique trigram count, while simultaneously aligning similarly to the gold responses, as reflected by BLEU scores. However, inform and success rates are slightly lower than those of PREV-gpt.
@@
To assess the impact of conditioning responses on explicit emotions, we conduct an in-house human evaluation using delexicalized Llama-2-based responses (Table 4). Our comparison involves baseline responses (SIMPLE), responses simply conditioned on user emotions through joint training (PREV), and these same PREV responses more explicitly conditioned via prompting (REFINE). We adopt a methodology inspired by nekvinda-dusek-2022-aargh (13) which involves side-by-side relative ranking, randomly selecting examples from each emotional category (neutral excluded). We create a collection of 60 evaluation examples, annotated by three distinct NLP practitioners.
@@ 4 Related work
Work relating to customer support dialogues focuses on the intersection of ED and TOD as these exchanges often stem from a poor user experience, making ED key. For instance, devillers-2003 (4) annotate emotions in a small corpus of 100 spoken call-center dialogues. zuters2020adaptive (27) also provide valuable insight, examining emotion intensity and frustration in customer support dialogues. However, the datasets utilized offer limited task-oriented annotations and lack explicit ties to a database, making it challenging to experiment with typical multi-domain TOD system baselines. wei2021target (24) take an interest in generating emotion-aware responses. In contrast to our work, they assume the user’s emotion is given and focus on a dataset of general social media post/response pairs, very different from the TODs to be found in MultiWOZ. feng-etal-2023-chatter (6) enhance emotion recognition models for chitchat and apply them to the complete EmoWOZ dataset. The full dataset contains an extra 1000 human-machine dialogues, not annotated with task information. Due to our specific approach, we focus solely on experimenting with the MultiWOZ portion of EmoWOZ and do not compare ED results with this study. Other research explores the use of emotionally-grounded user simulators for training dialogue policy models through reinforcement learning lin2023emous (11). In this setting, however, the system is not expected to predict the associated emotion label, in contrast to our endeavor.
@@
1.
SIMPLE: The address is [address] and the phone number is [phone].
(…) User: Can you tell me the address to the police station? I am in trouble and I need some help asap. (fearful)
PREV: The police station is located at [address].
REFINE: Oh no, I’m so sorry to hear that! The police station is located at [address].
2.
SIMPLE: I have [choice] options for you. [name] , [name] , and [name] . Do any of these sound good?
(…) User: It’s our anniversary and my wife loves british food. I want it to be special, so maybe something expensive. (excited)
PREV: I have [name] in the [area] that serves [food] food and is in the [pricerange] price range.
@@
3.
SIMPLE: The postcode for [name] is [postcode] and they do have free parking.
(…) User: You didn’t answer my question. I need the postcode for The Cambridge Belfry and I need to know if they have free parking or not. (abusive)
PREV: I’m sorry about that. The postcode for [name] is [postcode] . They do have free parking. Is there anything else I can help you with?
REFINE: I apologize, I did not see the question. I’m sorry about that. The postcode for [name] is [postcode] . They do have free parking. Is there anything else I can help you with?
4.
SIMPLE: Thank you for using our service.
(…) System: Is there anything else you need help with today?
PREV: Thank you for using our service.
User: That shall be all. Thanks! (satisfied)
REFINE: Yes you did! You’re very welcome! Thank you for using our service.
@@ 5 Conclusion
Our findings demonstrate that unifying ED and TOD modeling is overall beneficial for both tasks, producing improved results across ED and task-oriented metrics. Additionally, our experiments reveal that explicitly predicting emotion labels provides useful grounding for system responses, resulting in a heightened display of empathy especially when refining responses. In a nutshell, our findings underscore the advantages and ease of integrating ED into TOD modeling, compared with treating it as an independent task. We believe our contribution will stimulate TOD system developments which extend beyond the sole objective of task completion, fostering more user-focused innovations.
@@
We adapt the graphical user interface from nekvinda-dusek-2022-aargh (13) to fit our particular needs (Figure 2)
@@
User utterances and their associated emotion labels from EmoWOZ feng-etal-2022-emowoz (5) can be found in table 6. We direct readers to the paper for more details.
@@
Neutral
System: What sort of food would you like it to be?
@@
System: I have booked you clare hall
User: What? that’s not what i wanted.
Fearful, sad, disappointed
(Implicit expression)
@@
System: I’m afraid that’s all we’ve got.
Apologetic
User: I’m sorry I may have confused you. I would want to go from Petersborough to Cambridge on Saturday and arrive by 13:00, so anything that will get me there before that.
Abusive
System: (talking about something irrelevant)
User: *******, would ya stupid thing.
Excited, happy, anticipating
System: There are 3 restaurants on the west side that serve British food. Do you prefer moderate or expensive priced?
User: Expensive, please, we’re splurging for my mother’s birthday
Satisfied, liking, appreciative
System: Any time. I was happy to help.
User: You are doing a great job! That’s all I needed.
Table 6: Examples for each emotion label from EmoWOZ feng-etal-2022-emowoz (5).

### v3

# TowardsEmo.bib

@@ Abstract
In current task-oriented dialogue (TOD) systems, text-based user emotion detection (ED) is often overlooked or is typically treated as a separate and independent task, requiring additional training. In contrast, our work demonstrates that seamlessly unifying ED and TOD modeling brings about mutual benefits, and is therefore an alternative to be considered. Our method consists in augmenting SimpleToD, an end-to-end TOD system, by extending belief state tracking to include ED, relying on a single language model. We evaluate our approach using GPT-2 and Llama-2 on the EmoWOZ benchmark, a version of MultiWOZ annotated with emotions. Our results reveal a general increase in performance for ED and task results. Our findings also indicate that user emotions provide useful contextual conditioning for system responses, and can be leveraged to further refine responses in terms of empathy.
@@ 1 Introduction
Emotional user utterances frequently occur during interactions with dialogue systems reeves1996media. In open-domain dialogues, users directly share personal and emotional experiences li-etal-2017-dailydialog. In task-oriented dialogues (TODs), emotions are closely related to task progression. They usually become apparent as the task unfolds, and are contingent on the user’s expectations being met feng-etal-2022-emowoz. Detecting these emotions explicitly in either scenario is beneficial for several reasons: it can help with reviewing chat logs after the exchange concludes or with adjusting responses during the exchange to better align with the user’s emotional state rashkin-etal-2019-towards. In TODs, these more empathetic responses have demonstrated their effectiveness in compensating for system errors i-feel-you-2013, creating the impression of a more capable system.
Existing text-based approaches to ED in TODs either require a dedicated, specifically trained component feng-etal-2022-emowoz; feng-etal-2023-chatter or assume implicit ED. This assumption is due to systems being commonly trained to replicate human expert responses xu2017new; ham-etal-2020-end which inherently convey empathy, as needed, and therefore implicit ED. In contrast, we propose an approach that eliminates the need for training an additional component while explicitly modeling user emotions, applicable in settings where emotion annotations are available. Our method involves predicting user emotions within a sequence of task-related components, treating ED as an extension of belief state tracking. Our approach is end-to-end and relies on a single language model, optimized in a unified fashion. Specifically, we build upon SimpleTOD hosseini-asl_simple_2020, a popular end-to-end TOD approach trained on GPT-2 radford2019gpt2 (Sec. 2). Recognizing the success of large language models (LLMs) brown2020language, we also fine-tune a state-of-the-art LLM, Llama-2 touvron2023llama2, using LoRA hu2021lora, a parameter-efficient fine-tuning technique (Sec. 2).
Moreover, we show that when considering Llama-2, ED predictions can be readily used to more explicitly condition and refine system responses. Capitalizing on the fact that LoRA weights are optimized on top of Llama-2’s fixed parameters, we dynamically disable the LoRA weights after generating TOD predictions to harness the LLM’s in-context learning ability. We refine responses without further training, using a chain-of-thought prompting approach wei2022chain (Sec. 2).
We evaluate our method on the text-based EmoWOZ benchmark feng-etal-2022-emowoz, the only benchmark to provide emotion annotations on MultiWOZ budzianowski-etal-2018-multiwoz. The authors propose several competitive ED-only baselines, with ContextBERT feng-etal-2022-emowoz being the most successful and taking the entire dialogue history as input.
@@ 2 Method and Experimental Setup
This is the only text-based dataset offering emotion annotations for a large and popular corpus of 10,000 TODs. It focuses on 7 user emotions: neutral, excited, fearful, satisfied, dissatisfied, apologetic, and abusive. These follow the OCC model (Ortony, Clore and Collins) ortony_clore_collins_1988 and are based on 3 possible emotion elicitors which include the operator (or system), the user, and events. This emotion schema is highly specific to TODs and makes it possible to distinguish between emotions caused by the system and those provoked by external factors, independent of the system’s performance.
@@
Examples of user utterances and their associated emotion label can be found in Appendix 4. Dataset statistics are shown in Table 1. Certain emotions are notably rare and make this a challenging benchmark for ED in TODs.
@@
Our approach builds upon SimpleToD hosseini-asl_simple_2020, a popular end-to-end strategy that employs a single language model, similar to other recent end-to-end TOD systems ham-etal-2020-end; yang2021ubar. In training, the model is exposed to instances where Context (all preceding turns $C_{t}=[U_{0},S_{0},...,U_{t}]$), Belief State ($B_{t}$), Dialogue Acts ($A_{t}$) and System Response ($S_{t}$) are concatenated into a single text sequence. To ensure database adaptability, responses are additionally delexicalized. During inference, only the Context $C_{t}$ is passed as input, and the model is expected to generate the rest of the sequence. Moreover, the predicted belief state is used to query the database to retrieve values with which to replace delexicalized placeholders. This pipeline can be readily augmented, and previous studies have in fact done so with snippets of knowledge chen-etal-2022-ketod and chitchat sun-etal-2021-adding.
@@
The models used for our experiments include GPT-$2$https://huggingface.co/gpt2 radford2019gpt2, typically used with SimpleToD, and the more recent, state-of-the-art Llama-2-chat [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hftouvron2023llama2. This is an open-source LLM trained on data from publicly accessible sources which can undergo further fine-tuning or be prompted directly with instructions and/or few-shot examples. To adapt models to our task, we fully fine-tune GPT-2 and employ LoRA fine-tuning for Llama-2, preserving a significant portion of the pre-trained weights. This approach enhances regularization, generally leading to improved generalization compared to a fully fine-tuned alternative. Training details can be found in Appendix 2.
When refining responses, we focus on the PREV model’s predictions. Once these have been generated, we can disable the LoRA weights and, with no additional training, prompt the model using a few-shot, chain-of-thought wei2022chain approach (REFINE). This consists in creating demonstrations which include a reasoning step to prime the LLM’s output. Each few-shot demonstration consists in a) the dialogue Context, b) the emotion to be predicted, c) the response to be generated, d) a chain-of-thought reasoning step suggesting which behavior to display. Finally, it also contains e) an emotion-aware snippet that can be prepended to the previously generated response. Compared with refining the entire response, prompting the model to produce a prependable snippet mitigates task-related errors due to substantial modifications made by the LLM. We incorporate specific instructions into the prompt and construct one exemplar per emotion, picking examples from the training set. The demonstration snippets we design are meant to convey remorse, reassurance, appeasement, or enthusiasm, depending on the user’s emotion. Appendix 3 shows an illustrative example. When generating snippets, we filter those which are overly similar to the generated response to avoid redundancy, based on Levenshtein distance. Empirically, we find that filtering out snippets with a similarity ratio $\geq$ 50% is effective.
@@
For ED, we adhere to the EmoWOZ metrics, presenting F1 scores for each emotion alongside the macro and weighted averages excluding the neutral class. For end-to-end TOD, we employ the standard MultiWOZ metrics. The joint goal accuracy (JGA) reflects the proportion of turns where the predicted user constraints exactly match the reference ones. The inform rate evaluates the system’s capacity to provide the right type of entities from the database, given the user’s constraints, and the success rate assesses how effectively the system delivers requested attributes like phone numbers or booking references. For a more comprehensive understanding of these metrics, we direct readers to the MultiWOZ paper budzianowski-etal-2018-multiwoz. For response quality, we use the BLEU papineni-etal-2002-bleu score. For response diversity, we analyze the Conditional Bigram Entropy (CBE) and the count of unique trigrams to gauge the richness of vocabulary and phrasing nekvinda-dusek-2021-shades.
@@ 3 Results and Discussion
For Llama-2, the PREV variant shows a clearer benefit. This model performs best overall, except for the dissatisfied emotion. By examining user utterances with this particular emotion, we find that they appear highly contextual and generally lack explicit signs of dissatisfaction, a common trait when identifying emotions from text only multimodal_emo_rec. The regularization from LoRA in this case is a drawback compared to full fine-tuning, as annotations for this emotion are dataset-specific. However, for less common emotions characterized by more explicit expressions, such as fearful and abusive, the use of an LLM with LoRA proves advantageous. Nonetheless, opportunities for further improvements remain, particularly in the recognition of abusive and dissatisfied, two crucial emotional states.
@@
On the other hand, the EMO variant does not offer as much signal, yielding results similar to the baseline in each case. Compared to GPT-2, Llama-2 responses exhibit greater diversity, as indicated by the CBE score and the unique trigram count, while simultaneously aligning with the gold responses, as reflected by BLEU scores. However, inform and success rates are slightly lower than those of PREV-gpt.
@@
To assess the impact of conditioning responses on explicit emotions, we conduct an in-house human evaluation using delexicalized Llama-2-based responses (Table 4). Our comparison involves SIMPLE baseline responses, PREV responses conditioned on user emotions through joint training, and REFINE responses, which are PREV responses more explicitly conditioned via prompting. We adopt a methodology inspired by nekvinda-dusek-2022-aargh which involves side-by-side relative ranking, randomly selecting examples from each emotional category (neutral excluded). We create a collection of 60 evaluation examples, annotated by three distinct NLP practitioners.
@@ 4 Related work
Work relating to customer support dialogues has also focused on the intersection between ED and TOD, as these exchanges often stem from a poor user experience, making ED key. For instance, devillers-2003 annotate emotions in a small corpus of 100 spoken call-center dialogues. zuters2020adaptive also provide valuable insight, examining emotion intensity and frustration in customer support dialogues. However, the datasets utilized offer limited task-oriented annotations and lack explicit ties to a database, making it challenging to experiment with typical multi-domain TOD system baselines. wei2021target demonstrate a similar interest as ours by endeavoring to create emotion-aware responses. In contrast to our work however, they assume the user’s emotion is given. Additionally, they focus on a dataset of general social media post/response pairs, very different from the TODs to be found in MultiWOZ. feng-etal-2023-chatter enhance emotion recognition models for chitchat and apply them to the complete EmoWOZ dataset. The full dataset contains an extra 1000 human-machine dialogues, not annotated with task information. Due to our specific approach, we focus solely on experimenting with the MultiWOZ portion of EmoWOZ and do not compare ED results with this study. Other research explores the use of emotionally-grounded user simulators for training dialogue policy models through reinforcement learning lin2023emous. In this setting, however, the system is not expected to predict the associated emotion label, in contrast to our endeavor.
@@
1. SIMPLE: The address is [address] and the phone number is [phone].
(…) User: Can you tell me the address to the police station? I am in trouble and I need some help asap. (fearful) PREV: The police station is located at [address]. REFINE: Oh no, I’m so sorry to hear that! The police station is located at [address].
2. SIMPLE: I have [choice] options for you. [name] , [name] , and [name] . Do any of these sound good?
(…) User: It’s our anniversary and my wife loves british food. I want it to be special, so maybe something expensive. (excited) PREV: I have [name] in the [area] that serves [food] food and is in the [pricerange] price range.
@@
3. SIMPLE: The postcode for [name] is [postcode] and they do have free parking.
(…) User: You didn’t answer my question. I need the postcode for The Cambridge Belfry and I need to know if they have free parking or not. (abusive) PREV: I’m sorry about that. The postcode for [name] is [postcode] . They do have free parking. Is there anything else I can help you with? REFINE: I apologize, I did not see the question. I’m sorry about that. The postcode for [name] is [postcode] . They do have free parking. Is there anything else I can help you with?
4. SIMPLE: Thank you for using our service.
(…) System: Is there anything else you need help with today? PREV: Thank you for using our service.
User: That shall be all. Thanks! (satisfied) REFINE: Yes you did! You’re very welcome! Thank you for using our service.
@@ 5 Conclusion
Our findings demonstrate that unifying ED and TOD modeling is overall beneficial for both tasks, producing improved results across ED and task-oriented metrics. Additionally, our experiments reveal that explicitly predicting emotion labels provides useful grounding for system responses, resulting in a heightened display of empathy especially when refining responses. In a nutshell, our findings underscore the advantages and ease of integrating ED into TOD modeling, compared with treating it as an independent task. We believe our contribution will stimulate TOD system developments which extend beyond the sole objective of task completion, fostering more user-focused advancements benefiting various conversational settings campillos2020designing; kim2004effects.
6 Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 20XX-AD011014510 made by GENCI.
@@ 6 Acknowledgments
We adapt the graphical user interface from nekvinda-dusek-2022-aargh to fit our particular needs (Figure 2)
@@
EmoWOZ user utterances and their associated emotion labels can be found in Table 6. We direct readers to the paper for more details.
@@
Neutral System: What sort of food would you like it to be?
@@
Fearful, sad, disappointed User: Please help I’ve been robbed!!
@@
User: Actually, could you get me some information about museums in the area?
System: We actually have 23 restaurants in town! Is there a certain area you’ll be in?
User: I’m not looking for a restaurant, I want a museum !
Dissatisfied, disliking (Implicit expression)
@@
Apologetic System: There are several Italian Restaurants. What area are you interested in?
User: I am so sorry, I would rather have gastropub food, moderately priced please.
System: …There are 10 attractions in the east - do you care what sort?
Abusive User: Just do what I say! I don’t care what you choose! Give me the necessary information too.
Excited, happy, anticipating User: I’m traveling to Cambridge and looking forward to their restaurants ! I need a place to stay while I’m there.
Satisfied, liking, appreciative System: Is there anything else I can assist you with?
User: No, thanks you’ve been great!
Table 6: Examples for each emotion label from EmoWOZ feng-etal-2022-emowoz.
[…]